Fill in the blanks in the table below. Assume that the MPC is constant over everyone in the economy.

MPC Spending Multiplier Change in Government Spending Change in Income
10 100
2.5 -500
0.5 225
0.2 100
Answer #1
MPC Spending Multiplier= (1/1-MPC) Change in government spending= (Change in income/Spending multiplier) Change in income = (Multiplier)(Change in government spending)
0.9 10 100 1000
0.6 2.5 -200 -500
0.5 2 225 450
0.2 1.25 80 100

Spending multiplier = 1/1-MPC

10 = 1/1-MPC

MPC = 1-0.1

MPC = 0.9